KpopAsianDramaFanYT:

Please help with these questions!!! Q1: Show all work to factor x^4 − 5x^2 + 4 completely. Q2: Show all work to solve 3x^2 − 5x − 2 = 0. Q3: Two functions are given below: f(x) and h(x). State the axis of symmetry for each function and explain how to find it. f(x) h(x) f(x) = −4(x − 8)^2 + 3 graph of negative 1 times the quantity of x minus 3 squared, plus 2 Q4: The functions f(x) = (x + 1)^2 − 2 and g(x) = −(x − 2)^2 + 1 have been rewritten using the completing-the-square method. Apply your knowledge of functions in vertex form to determine if the vertex for each function is a minimum or a maximum and explain your reasoning.

KpopAsianDramaFanYT:

Cancel the first 2 Q's

KpopAsianDramaFanYT:

only Q 3 and 4 i need help on

jhonyy9:

x^4 -5x^2 +4 = x^2 *x^2 - 4x^2 -x^2 +(-1)(-4) = x^2 ( x^2 -4) -(x^2 -4) = (x^2 -4)(x^2 -1)

KpopAsianDramaFanYT:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @jhonyy9 x^4 -5x^2 +4 = x^2 *x^2 - 4x^2 -x^2 +(-1)(-4) = x^2 ( x^2 -4) -(x^2 -4) = (x^2 -4)(x^2 -1) \(\color{#0cbb34}{\text{End of Quote}}\) i said that I didn't need help on q1 and 2 only q 3 and 4

YECR:

Sorry Bro but we are only allowed to Guide you To the answer We cant directly give you The Answer cause that's direct Answer.

kaleighh533:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @YECR Sorry Bro but we are only allowed to Guide you To the answer We cant directly give you The Answer cause that's direct Answer. \(\color{#0cbb34}{\text{End of Quote}}\) YECR right if anyone gets caught giving direct answer's they have to be warned so we cannot give the the direct answers only lead u to them to were u understand.

YECR:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kaleighh533 \(\color{#0cbb34}{\text{Originally Posted by}}\) @YECR Sorry Bro but we are only allowed to Guide you To the answer We cant directly give you The Answer cause that's direct Answer. \(\color{#0cbb34}{\text{End of Quote}}\) YECR right if anyone gets caught giving direct answer's they have to be warned so we cannot give the the direct answers only lead u to them to were u understand. \(\color{#0cbb34}{\text{End of Quote}}\) Exactly

hotboy3245657868760:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @kaleighh533 \(\color{#0cbb34}{\text{Originally Posted by}}\) @YECR Sorry Bro but we are only allowed to Guide you To the answer We cant directly give you The Answer cause that's direct Answer. \(\color{#0cbb34}{\text{End of Quote}}\) YECR right if anyone gets caught giving direct answer's they have to be warned so we cannot give the the direct answers only lead u to them to were u understand. \(\color{#0cbb34}{\text{End of Quote}}\) YERC and kaleighh533 are right

YECR:

@BabyGurl2006

BabyGurl2006:

if u figured out q1 and q2...im sure u know things about the functions

BabyGurl2006:

(hint hint)

kaleighh533:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @BabyGurl2006 if u figured out q1 and q2...im sure u know things about the functions \(\color{#0cbb34}{\text{End of Quote}}\) thats also true bcs question 2 kinda looks like question 3 but maybe they just didnt understand it

KpopAsianDramaFanYT:

I just need help with Q3 & Q4

KpopAsianDramaFanYT:

mostly Q4